OpenStudy (anonymous):

I'v done this kind of stuff before and I just don't remember how Solve for U T=3U/E

Mehek (mehek14):

\(\bf{T=\dfrac{3U}{E}\\ or \\T=3\dfrac{U}{E}}\)

OpenStudy (anonymous):

the first equation you have there

Mehek (mehek14):

first thing, multiply E on both sides \(T•E=3U\)

Mehek (mehek14):

next, divide 3 on both sides \(\bf{\dfrac{T•E}{U}=3}\)

OpenStudy (anonymous):

ok i see thank you

Mehek (mehek14):

basically, your goal is to move everything to the opposite side

Mehek (mehek14):

yw and welcome to OpenStudy \(\ddot\smile\)

OpenStudy (anonymous):

thank you